OpenStudy (anonymous):

find the missing term of the arithmetic sequence 22 _, 34

OpenStudy (anonymous):

average

OpenStudy (anonymous):

what is half way between 22 and 34?

OpenStudy (triciaal):

remember for an AP the terms of the sequence differ by the same amount 34 - ? = ? -22 34 + 22 = 2*? ? = 56/2 = 28 to check 34-28 = 28 - 22 6 + 6 therefore the AP 22, 28, 34

OpenStudy (triciaal):

do you have any questions about this?